OpenStudy (adamaero):

Integral from infinity to negative infinity of Asin^2(x)+Bx^(-2) dx (all equal to one)

OpenStudy (adamaero):

\[1 = \int\limits_{-\infty}^{\infty}Asin^2(\alpha x)+Bx^{2}dx\] raw text: 1 = int_{-infty}^{infty}Asin^2(alpha x)+Bx^{2}dx

OpenStudy (irishboy123):

is it odd or even?

OpenStudy (adamaero):

Sin(x) is odd: f(–x) = –f(x) A squared function is even.

OpenStudy (loser66):

On the question box, you said Bx^(-2) On comment box, you said Bx^2 Which one is the correct one?

OpenStudy (adamaero):

Are you getting at something about the four conditions that need to be satisfied for a wave function? (1) Cont. (2) No corners or cusps (3) Finite area (4) squared integral (normalizable)

OpenStudy (irishboy123):

yes!!

OpenStudy (loser66):

And we have to prove it?

OpenStudy (adamaero):

I need to find A & B.

OpenStudy (adamaero):

\[1 = \int\limits_{-\infty}^{\infty}Asin^2(\alpha x)+Bx^{-2}dx\] raw text: 1 = int_{-infty}^{infty}Asin^2(alpha x)+Bx^{-2}dx

OpenStudy (irishboy123):

if this is a wave function, you also have some IV's are you presenting all of the question?!

OpenStudy (adamaero):

A wavefunction for a particle confined in some region of space is given by the expression: φ(x)=Asin^2(αx)+Bx^(-2) What are the dimensions (i.e., units) of A and B?

OpenStudy (irishboy123):

lol!!

OpenStudy (irishboy123):

i think the wave eq is dimensionless and so is A B is oppo of 1/m^2 that's not yer question, though?!?!

OpenStudy (adamaero):

Nope, that's not my question. I thought that's what I was suppose to do. Why are they opposite of that? And what is m?

OpenStudy (irishboy123):

ask something similar in the physics forum :-)

OpenStudy (adamaero):

That's why I tried to turn it into a math problem (very few people available for the other subjects)

OpenStudy (adamaero):

\[1 = \int\limits_{-\infty}^{\infty}Asin^2(\alpha x)+Bx^{-2}dx\] Integrating something basically multiplies it by the dimension of integration. So integrating the modulus multiplies the result by meters: \[1 = \int\limits_{-\infty}^{\infty}1/(Asin^2(\alpha x)+Bx^{-2})dx\] Right? the modulus??